# Target Tracking via LiDAR-RADAR Sensor Fusion for Autonomous Racing

### v2

@@ III-D EKF Measurement Function
The EKF measurement function $\mathbf{H}(X)$ is time-varying and size-varying, as the shape of the measurement depends on the type of sensor employed: it can be $Y_{L}=[x,y]\in\mathbb{R}^{2}$ or $Y_{R}=[x,y,R_{r}]\in\mathbb{R}^{3}$. For both LiDAR and RADAR measurements, we apply the transformation from local (EGO) to global (earth-fixed) reference frame for the ${x,y}$ position measurements, in order to have them in the same reference frame as the EKF model.
@@
The LiDAR measurement function is linear, as after the reference frame transformation the LiDAR directly measures the first two states, giving $Y(k)=H_{L}\cdot X(k)$, from which derives:
@@
The RADAR measurement function instead is non-linear and time-varying as the Range Rate represents the relative EGO-target velocity expressed in the EGO reference frame and projected on the Line-of-Sight defined by the bearing angle $\alpha$, as Figure 4 shows. Therefore, the RADAR measurement function depends not only from the target states but also on the EGO velocity and heading, which are assumed to be known.
Figure 4: Representation of the components of the Range Rate measurements: Line-of-Sight angle $\alpha$, etc…
Given the knowledge of the EGO velocity and heading ${v_{E}},\theta_{E}$, the Range Rate equation is:
@@
III-E EKF Prediction and Data Association
The first step of every MTT cycle is EKF Prediction, in which every track is subject to motion prediction according to Equation (3). The yaw rate input $u$ is computed according to (4) for every track and then used to update the state estimation.
During the prediction step, we apply (1) in order to compute both the predicted state $\hat{X}(k+1)$ and its corresponding predicted output $\hat{Y}$. We also update the current state and output error covariance estimation according to:
$\hat{X}(k)$ $=\mathbf{F}(X_{k-1},u_{k})$ (8)
$\hat{Y}(k)$ $=\mathbf{H}(\hat{X}_{k})$
$A(k)$ $=\left.\frac{d\mathbf{F}}{dx}\right|_{\hat{X}_{k}}$
$P(k)$ $=A_{k}P_{k-1}A_{k}^{T}+Q$
$S(k)$ $=H_{L,R}P_{k}H_{L,R}^{T}+R$
where $A$ represents the state linearization around the current prediction, $P$ the state error covariance estimate and $S$ the output error covariance estimate, while $Q$ and $R$ represent the model and output noise covariance matrices.
@@ III-E EKF Prediction and Data Association
We evaluate the Mahalanobis Distance for every track-measure pair, building a cost matrix. We impose an upper Gating Threshold on the computed Mahalanobis distance in order to inhibit the association of measures that are too far away from the corresponding track, and then we use the Munkres Algorithm to find the track-measurement assignment with the lowest overall cost.
The use of the Mahalanobis distance for data association allows the data association process to encompass information about the current track state estimation confidence: a track that has not been corrected for long will have a larger association region to reflect the larger uncertainty associated to the current vehicle pose estimate.
@@
After the Data Association step, measurements associated to a track will be used to correct the EKF state prediction, while those measurements that have not been associated to an existing track will be used to spawn a new track.
III-F Track Management
@@
Figure 5: Track Management Finite State Machine states and transition
After the Data Association step, we know which tracks have been associated with a measure in the last iteration and which have not. This information is accumulated and used to enable the transitions of the Track Management Finite State Machine represented in Figure 5.
@@
For every track, we define a FSM with four states: Tentative, Accepted, Confirmed and Terminated. Every track FSM is initialized in the Tentative state, with the Terminated state being the terminal state, after which the track is deleted. Only Confirmed tracks represent the output of the Target Tracking algorithm, while Accepted tracks represent an intermediate state.
@@
The state transition function of the FSM depends on the so-called M/N Logic, meaning that state transitions are enabled if the corresponding track has received at least N measurements in the last M MTT cycles. We define two upper thresholds, $TH_{acc}$ and $TH_{conf}>TH_{acc}$ for the transition from Tentative to Accepted and from Accepted to Confirmed respectively. A lower threshold, $TH_{elim}<TH_{acc}$ is used to transition tracks from either the Accepted or Confirmed states to the Terminated state if not met, effectively eliminating them. The elimination of Tentative tracks happens if, once M measurements have been received, the track has not been Accepted, to quickly eliminate false positives.
III-G EKF Correction
@@
The last step of the MTT cycle is the actual EKF correction, in which we use the measurement associated to every track (except Terminated tracks) to correct the estimation of its dynamical state using the EKF Update equations:
$E(k)$ $=(Y_{k}-\hat{Y}_{k})$ (10)
$L(k)$ $=P_{k}\cdot H_{L,R}^{T}(k)\cdot S^{-1}_{k}$
$X(k)$ $=X_{k-1}+L\cdot E_{k}$
where $E$ is the measurement error or innovation of associated measurement $Y$ and $L$ is the Kalman Gain.
@@
This EKF correction is not applied to the first associated measurement of a track, which determines the track creation and is initialized with $v=0,\theta=0$, nor for the second associated measurement. To speed up the EKF convergence, we perform the Two Step Initialization procedure described in [14]. During this procedure we compute the initialization velocity and heading as:
$v_{2}$ $=\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}}$ (11)
$\theta_{2}$ $=\arctan{((y_{2}-y_{1})/(x_{2}-x_{1}))}$
@@
This procedure quickly initializes the track’s velocity and heading to values that, although noisy, enable faster convergence to the true values compared to a default zero initialization.
@@ IV EXPERIMENTAL RESULTS
The vehicle used for the experimental data collection and algorithm validation is a Dallara AV21 minerVa. It is a single-seater autonomous race car equipped with three Luminar Hydra LiDAR sensors, positioned in order to have 360° coverage around the vehicle (although with some minor blind spots), and a single forward-facing APTIV ESR 2.5 RADAR sensor. The vehicles are equipped with two NovAtel PwrPak7D RTK INS/GNSS receivers which act both as the EGO vehicle localization source for the algorithm operations as well as a Ground Truth source for performance evaluation.
@@
Figure 6: Map and acceleration profile of the two racing circuits used for this work: on the left, the Las Vegas Motor Speedway, on the right the Autodromo Nazionale di Monza. Distances are expressed in meters.
We evaluate the algorithm performance on two datasets: the first is an experimental dataset taken at the Las Vegas Motor Speedway during the 2023 IAC event. It consists of 15 overtaking maneuvers captured both by the overtaking and overtaken vehicle perspective at speeds ranging from 160 to 275 km/h. Due to the nature of the oval racetrack, as well as the racing regulations which imposed specific speed brackets for the overtaken vehicle, this dataset is characterized by long sequences at almost constant EGO and opponent speed.
@@
To overcome this limitation, we generated the second dataset using the experimental data acquired during the PoliMOVE qualifying lap at the 2023 IAC event at the Autodromo Nazionale di Monza, by simulating a chasing vehicle following a leader using the recorded trajectories of the real vehicle. We added noise to the ideal sensor measurement with a zero-mean Gaussian model, following the noise characterization of the real sensor data w.r.t. the Ground Truth. Figure 6 shows the different acceleration profile of the two datasets, showing the highly variable vehicle speed of the Monza Racetrack.
Figure 7: Experimental measurement and tracking errors w.r.t. the Ground Truth on the LVMS Dataset. On the left, 1-sigma ellipses for the EGO x and y coordinates. On the right, histograms of speed and heading error.
We evaluate the tracking performance by measuring the absolute difference between the filtered track states and the ground truth data acquired from the RTK GNSS. To distinguish between forward (longitudinal) and lateral offsets, we transform both tracked and ground-truth states into an EGO-centered coordinate system, computing errors in the local $x$ and $y$. Moreover, we utilize identical EKF parameters and noise covariances for both the real-world LVMS dataset and the simulated Monza dataset. This uniform setup allows us to assess how well the algorithm generalizes to different track geometries and acceleration profiles, thus confirming its robustness in diverse racing scenarios.
Figures 7 and 8 show the tracking error of the algorithm on both datasets. We can see that the tracking filter is successfully capable of reducing the estimation error of the measurable states, as well as provding a reasonable estimation of the non-measurable states. The real LVMS data present a bias on the x direction, which derives from sensor mis-calibration, and different mean values between LiDAR and RADAR on the y direction, which are compensated by the EKF. Although in principle the CVTR model is more suited to low acceleration circuits like LVMS, it still manages to provide a consistent tracking and noise reduction on a circuit with a more aggressive longitudinal and lateral acceleration profile. Note that the largest error component is in the lateral direction also on the simulated data, although its acceleration profile was supposed to stress more the longitudinal dynamics estimation.
@@ V CONCLUSIONS
In this paper, we present a latency-aware EKF-based multi-target tracking framework that integrates LiDAR and RADAR measurements, explicitly incorporating track curvature in the prediction step and RADAR range-rate information in the correction step. By managing Out-Of-Sequence Measurements through state and measurement reprocessing, the proposed can effectively handle the different delays deriving from multi-modal fusion.
Figure 8: Experimental measurement and tracking errors w.r.t. the Ground Truth on the Monza simulated dataset. On the left, 1-sigma error ellipses for the local x and y coordinates. On the right, histograms of speed and heading error.
Experimental validation on real-world data from the Las Vegas Motor Speedway at speeds up to 275 km/h, as well as a simulated dataset generated from racing laps at the Autodromo Nazionale di Monza, demonstrated robust and reliable estimation performance under widely varying conditions and acceleration profiles.

### v3

@@ III-C EKF Measurement Function
The EKF measurement function $\mathbf{H}(X)$ is time-varying and size-varying, as the shape of the measurement depends on the type of sensor employed: it can be $Y_{L}=[x,y]\in\mathbb{R}^{2}$ or $Y_{R}=[x,y,R_{r}]\in\mathbb{R}^{3}$. For both LiDAR and RADAR measurements, we apply the transformation from local (ego) to global (earth-fixed) reference frame for the x,y{x,y} position measurements, in order to have them in the same reference frame as the EKF model.
@@
The LiDAR measurement function is linear, as after the reference frame transformation the LiDAR directly measures the first two states, giving $Y_{k}=H_{L}\cdot X_{k}$, from which derives:
@@
The RADAR measurement function instead is non-linear and time-varying as the range-rate represents the relative ego-target velocity expressed in the ego reference frame and projected on the Line-of-Sight defined by the bearing angle $\alpha$, as Figure 5 shows. Therefore, the RADAR measurement function depends not only on the target states but also on the ego velocity and heading, which are assumed to be known. Given the knowledge of the ego velocity and heading ${v_{e}},\theta_{e}$, in the hypothesis of small ego sideslip angle, the range-rate equation is
@@
Figure 5: Representation of the components of the RADAR Doppler measurement function: the range-rate is the projection of the relative velocity $v_{rel}$ over the Line-of-Sight angle $\alpha$.
III-D Data Association and Track Management
During the MTT prediction step, we apply (1) to compute both the predicted state $\hat{X}_{k+1}$ and its corresponding predicted output $\hat{Y}$ for every track, using (5) or (7) depending on the sensor type. We also update the current state and output error covariance estimation according to:
$\hat{X}_{k}$ $=\mathbf{F}(X_{k-1},u_{k})$ (8)
$\hat{Y}_{k}$ $=\mathbf{H}(\hat{X}_{k})$
$F_{k}$ $=\left.\frac{d\mathbf{F}}{dx}\right|_{\hat{X}_{k}}$
$P_{k}$ $=F_{k}\cdot P_{k-1}\cdot F_{k}^{T}+Q$
$S_{k}$ $=H_{k}\cdot P_{k}\cdot H_{k}^{T}+R$
where $F$ represents the state linearization around the current prediction, $P$ the state error covariance estimate and $S$ the output error covariance estimate, while $Q$ and $R$ represent the model and output noise covariance matrices.
@@ III-D Data Association and Track Management
We evaluate the Mahalanobis Distance for every track-measurement pair, building a cost matrix. We impose an upper gating threshold on the computed Mahalanobis distance in order to inhibit the association of measurements that are too far away from the corresponding track, and then we use the Munkres Algorithm to find the track-measurement assignment with the lowest overall cost. The use of the Mahalanobis distance for data association allows the data association process to encompass information about the current track state estimation confidence: a track that has not been corrected for long will have a larger association region to reflect the larger uncertainty associated with the EKF position estimate.
@@
After the Data Association step, measurements associated with a track will be used to correct the EKF state prediction, while every measurement that has not been associated with an existing track will be used to spawn a new track.
@@
Figure 6: Track Management Finite State Machine states and transitions
@@
For every new track, we instantiate a Finite-State Machine (FSM), shown in Figure 6, with four states: Tentative, Accepted, Confirmed and Terminated. Every track FSM is initialized in the Tentative state, with the Terminated state being the terminal state. Only Confirmed tracks represent the output of the Target Tracking algorithm, while Accepted tracks represent an intermediate state.
@@
The state transition function of the FSM depends on the so-called M/N Logic, meaning that state transitions are enabled if the corresponding track has received at least N measurements in the last M MTT cycles. We define two upper thresholds, $TH_{acc}$ and $TH_{conf}>TH_{acc}$ for the transition from Tentative to Accepted and from Accepted to Confirmed respectively. A lower threshold, $TH_{elim}<TH_{acc}$, is used to transition tracks from either the Accepted or Confirmed states to the Terminated state if not met, effectively eliminating them. The elimination of Tentative tracks happens if, once M measurements have been received, the track has not been Accepted.
III-E EKF Correction and Initialization
@@
The last step of the MTT cycle is the actual EKF correction, in which we use the measurement associated with every track (except Terminated tracks) to correct the estimation of its dynamical state using the EKF update equations as in
$E_{k}$ $=(Y_{k}-\hat{Y}_{k})$ (10)
$L_{k}$ $=P_{k}\cdot H_{k}^{T}\cdot S^{-1}_{k}$
$X_{k}$ $=X_{k-1}+L_{k}\cdot E_{k}$
where $E$ is the measurement error or innovation of the associated measurement $Y$ and $L$ is the Kalman Gain.
@@
This EKF correction is not applied to the first associated measurement of a track, which determines the track creation and is initialized with $v=0,\theta=0$, nor for the second associated measurement: to speed up the EKF convergence, we perform the two-step initialization procedure described in [bar2004estimation]. During this procedure we compute the initialization velocity and heading as:
$v$ $=\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}}$ (11)
$\theta$ $=\operatorname{atan2}{(y_{2}-y_{1},x_{2}-x_{1})}$
@@
This procedure quickly initializes the track’s velocity and heading to values that, although noisy, enable faster convergence to the true values compared to the zero initialization.
@@ IV EXPERIMENTAL RESULTS
Figure 7: Map and acceleration profile of the two racing circuits used for this work: on the left, the Las Vegas Motor Speedway, on the right the Autodromo Nazionale Monza. Distances are expressed in meters.
The vehicle used for the experimental data collection and algorithm validation is the Dallara AV-21 shown in Figure 1. It is a single-seater autonomous race car equipped with three Luminar Hydra LiDAR sensors, positioned to provide almost 360° coverage around the vehicle, and a single forward-facing APTIV ESR 2.5 RADAR sensor. The vehicle is equipped with two NovAtel PwrPak7D RTK INS/GNSS receivers which act both as the vehicle localization source for the algorithm operations as well as the opponent ground truth source, as all IAC cars follow the same specification.
@@
We evaluate the algorithm’s performance on two datasets: the first is an experimental dataset taken at the Las Vegas Motor Speedway during the 2023 IAC event, consisting of 15 overtaking maneuvers captured from both the overtaking and overtaken vehicles’ perspectives, up to 160 to 275 km/h. Due to the oval racetrack, as well as the regulationsimposing specific speed brackets for the overtaken vehicle, this dataset is characterized by long sequences at almost constant speeds. We generated a second dataset using the experimental data acquired during the PoliMOVE qualifying lap at the 2023 IAC event at the Autodromo Nazionale Monza, by simulating a chasing vehicle following a leader using the recorded trajectories of the real vehicle. We added noise to the ideal sensor measurements with a zero-mean Gaussian model, following the noise characterization of the real sensor data with respect to the ground truth. Figure 7 shows the different acceleration profile of the two datasets, showing the highly variable vehicle speed on the Monza track.
The online ROS2 implementation yields an average computational time under 1 ms, which is negligible with respect to the LiDAR processing latency of 26ms [cellina2025lidar] as well as the estimated RADAR latency of 100ms. The latency compensation via the final EKF state prediction compensates the end-to-end latency of the VDT pipeline.
Figure 8: 2D position error ablation study on the LVMS dataset
Figure 9: 2D position error ablation study on the Simulated Monza dataset
We compute the 2D position error to evaluate the performance of the tracking algorithm, and then we perform an ablation study by disabling each component one at a time in order to understand their contribution to the overall performance. Figure 8 displays the distribution of the raw measurements and of the Tracking state estimation 2D position error computed with respect to the GNSS ground truth for the LVMS dataset. The figure shows how the LiDAR has an overall lower 2D position estimation error compared with RADAR, although with higher variance, whereas sensor-fusion tracking attains a lower average error, evidencing the filter’s noise-reduction effect. The deactivation of each sensor increases the median, with the RADAR-only tracking being noisier than the LiDAR-only tracking, reflecting the larger position uncertainty of the latter. The deactivation of the Yaw Rate Exogenous Input, meaning the estimation of $\omega$ as the fifth EKF state using the full-state CVTR model, does not significantly reduce the filter’s performance, while the disabling of OOSM management, by simply discarding OOSMs when they are present, greatly degrades tracking accuracy, as half of the measurements are discarded.
Figure 9 shows the result of the 2D position error ablation study for the simulated Monza dataset. While the full tracking error precision, as well as the trend in the ablation study, is comparable to the experimental data, the contribution of the Virtual Yaw Rate Exogenous Input to the reduction of the state estimation error is greater in the Monza dataset than it was in the LVMS dataset. This can be explained by considering the higher vehicle yaw rate dynamics in a road course circuit presenting more abrupt direction changes compared with an oval raceway. In general, the positive effect of direct state measurement or input (velocity via RADAR Doppler range-rate measurements, and yaw rate via the Exogenous Input) is more distinguishable in the simulated Monza dataset, due to the more aggressive acceleration profile.
@@ V CONCLUSIONS
In this paper, we present a latency-aware, EKF-Based Multi-Target Tracking algorithm that integrates LiDAR and RADAR measurements, explicitly incorporating track curvature in the prediction step and RADAR range-rate information in the measurement model. By managing Out-Of-Sequence Measurements through state and measurement reprocessing, the proposed method can effectively handle the different delays arising from multi-modal fusion. Experimental validation on real-world data from the Las Vegas Motor Speedway, as well as the simulated Monza dataset, demonstrates robust and reliable estimation performance under widely varying conditions and acceleration profiles. Ablation studies on both datasets show the contribution of each module to reducing the estimation error.